OpenStudy (anonymous):

(3m^4)/(2m^2n) divided by (15m^3n^2)/(16m)

OpenStudy (anonymous):

do you know how to solve this much simpler problem ? 3/2 divided by 15/16

OpenStudy (anonymous):

yeah 8/5

OpenStudy (anonymous):

how did you do it ?

OpenStudy (anonymous):

show all steps

OpenStudy (anonymous):

3/2 divided by 15/16; 3/2 16/15; cross out 3 and get one then cross out 15 and get 5; cross out to and get one then cross out 16 and get 8. 8/5

OpenStudy (anonymous):

do you know this rule ? \[\frac{(\frac{a}{b})}{(\frac{c}{d})}=\frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait, maybe

OpenStudy (anonymous):

it's just a way to make a big division problem look less scary

OpenStudy (anonymous):

ok does it mean turn the division into mutipication?

OpenStudy (anonymous):

on the left side we have something very scary (4 levels high), and on the right side we have the same thing, but it looks like a normal fraction.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yes you turn a division into multiplication. instead of dividing, you multiply by the reciprocal. so you know this method ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

now, do you know how to solve this: \[\frac{5^5}{5^2}\]

OpenStudy (anonymous):

1\[^{3}\]

OpenStudy (anonymous):

1^3

OpenStudy (anonymous):

\[\frac{5^5}{5^2}=\frac{5^3}{5^0}=5^3=5*5*5=125\]

OpenStudy (anonymous):

see what happened there ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Look carefully at what happens to the powers of 5

OpenStudy (anonymous):

oh, I get it

OpenStudy (anonymous):

when we divide

OpenStudy (anonymous):

try this one: \[\frac{5^3}{5^5}=?\]

OpenStudy (anonymous):

1/^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The number 5 should appear in your answer.

OpenStudy (anonymous):

5^0/5^2?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how about this: 5^3 * 5^2 = ?

OpenStudy (anonymous):

25^5

OpenStudy (anonymous):

5^3 * 5^2 = 5^(3+2) = 5^5

OpenStudy (anonymous):

but 5*5 is 25

OpenStudy (anonymous):

5*5=5^2=25 In multiplcation we add the powers of 5: 5*5=5^1 * 5^1 = 5^(1+1) = 5^2 In division we substract the powers of 5: 5^3 / 5^5 = 5^(3-5) = 5^(-2) = 1/5^2

OpenStudy (anonymous):

If any of the above doesnt make sense let me know :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now try these: m^2 * m^3 = ? m^3/m^4 =? m^5/m^3=?

OpenStudy (anonymous):

so whats the answer to the problem (3m^4)/(2m^2n) divided by (15m^3n^2)/(16m)

OpenStudy (anonymous):

do you just want the answer, or to be able to answer the question yourself easily ?

OpenStudy (anonymous):

I wanna know how to do it to

OpenStudy (anonymous):

You already solved for the number in the beginning remember ? you got the 8/5 part of the answer.

OpenStudy (anonymous):

now we need to focus on the m's and the n's .

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for example the m's what is: m^4/m^2 = ?

OpenStudy (anonymous):

m^2

OpenStudy (anonymous):

good. now what is m^3/m=?

OpenStudy (anonymous):

m^2

OpenStudy (anonymous):

so what is: ( m^4/m^2) / (m^3/m) =?

OpenStudy (anonymous):

m/m

OpenStudy (anonymous):

its: (m^2)/(m^2) = 1

OpenStudy (anonymous):

which means all the m's cancel out, and your final answer will have no m in it!

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

does that make sense ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

so lastly we need to look at the n's

OpenStudy (anonymous):

n^3

OpenStudy (anonymous):

we have: (1/n) / (n^2/1) = (1/n) * (1/n^2) = 1/n^3

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

so now we need to get the final answer, which is a product of the numeric part of the answer times the m's part of the answer times the n's part of the answer. that's how we get to: \[\frac{8}{5n^3}\]

OpenStudy (anonymous):

to summarize: \[\frac{(\frac{3m^4}{2m^2n})}{(\frac{15m^3n^2}{16m})}\]

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

\[=\frac{3m^4}{2m^2n}*\frac{16m}{15m^3n^2}=\frac{3*16*m ^{4+1}}{2*15*m^{2+3}n^{1+2}}=\frac{8m^5}{5m^5n^3}=\frac{8m^0}{5n^3}=\frac{8}{5n^3}\]

OpenStudy (anonymous):

so we only did two things. First we converted a big division problem into a smaller multiplication problem. Then, we just did some addition and subtraction on the powers of m and powers of n.